### v3

# ProtoTS: Learning Hierarchical Prototypes for Explainable Time Series Forecasting

Abstract
While deep learning has achieved impressive performance in time series forecasting, it becomes increasingly crucial to understand its decision-making process for building trust in high-stakes scenarios. Existing interpretable models often provide only local and partial explanations, lacking the capability to reveal how heterogeneous and interacting input variables jointly shape the overall temporal patterns in the forecast curve. We propose ProtoTS, a novel interpretable forecasting framework that achieves both high accuracy and transparent decision-making through modeling prototypical temporal patterns. ProtoTS computes instance-prototype similarity based on a denoised representation that preserves abundant heterogeneous information. The prototypes are organized hierarchically to capture global temporal patterns with coarse prototypes while capturing finer-grained local variations with detailed prototypes, enabling expert steering and multi-level interpretability. Experiments on multiple realistic benchmarks, including a newly released LOF dataset, show that ProtoTS not only exceeds existing methods in forecast accuracy but also delivers expert-steerable interpretations for better model understanding and decision support.
1 Introduction
Time series forecasting has been widely applied in high-stakes scenarios such as load forecasting (jiang2024interpretable; yang2023interactive), energy management (deb2017review; weron2014electricity), weather prediction (angryk2020multivariate; karevan2020transductive), all of which involve considerable financial impacts. In these applications, while achieving high forecast accuracy is crucial, understanding why and how the model makes specific predictions is equally important. It aids in preventing substantial financial losses and building the trust necessary (rojat2021explainable).
A range of explainable time series forecasting methods have been developed to simultaneously ensure interpretability and good predictive performance (oreshkin2019n; lim2021temporal; zhao2024disentangled; lin2024cyclenet). However, their overall interpretability and potential for further performance improvement are limited, since they mainly provide local, partial explanations for both the output and input sides:
  • C1: For the output side, existing methods (lim2021temporal; zhao2024disentangled) mainly explain the prediction at individual time steps, lacking the ability to help users quickly interpret the reasons behind the overall trend in the forecast curve. For example, they fail to explain why three peaks occur in the predicted electricity load curve at the noon, afternoon, and night, respectively, and why the three peaks gradually descend (Figure 1(a1)). Understanding such overall temporal patterns is of great importance in real-world scenarios. For example, power system experts need to ensure demand peaks are identified correctly, so that they can make dispatch decisions (6721569), such as whether additional electricity should be purchased from external sources. Failing to provide such analyses limits interpretability and prevents experts from steering the model to improve its accuracy.
  • C2: For the input side, existing explanations are often limited to certain types of variables, such as focusing solely on endogenous variables (lin2024cyclenet). However, as shown in Figure 1(a), there are many different types of input variables, and understanding how their interactions impact the temporal patterns is critical. For example, the typical temporal pattern of extreme summer conditions can only be observed when both temperature and seasonal variables are considered together.
Figure 1: (a) An example of time series forecasting with numerous heterogeneous variables, where exogenous variables (e.g., temperature, holiday) influence the evolution of endogenous variables (e.g., electric load). (b) Illustration of prototypical explanation method: a set of learned prototypes provides a user-friendly global overview of typical temporal patterns. For each instance, model computes its similarity to all prototypes to form a prediction, enabling detailed local interpretation.
To address these challenges, we propose a Prototypical Time Series Forecasting framework (ProtoTS) that effectively models the overall temporal patterns (C1) and enables prototypes to capture the interactions among inputs (C2). As shown in Figure 1(b), each of our prototypes corresponds to a typical temporal pattern. For example, the curve of the Spring Festival prototype [^1] The name “Spring Festival” is now manually summarized based on relevant covariates and temporal patterns, though it can also be generated using large language models. indicates that there are peaks in the morning and evening during the Spring Festival period. This pattern occurs at specific combinations of covariates (i.e., when the inputs indicate both a holiday and the first month of the lunar calendar), enabling it to explain both the formation of the overall temporal pattern and how it is influenced by the combination of multiple input features (C2). These few prototypes provide a global explanation by summarizing typical patterns across the entire dataset, thereby supporting expert intervention (e.g., refining the temporal pattern of the prototype, Sec. 4.4). At the same time, the model improves accuracy by effectively modeling the local interactions of covariates.
To construct an effective end-to-end framework, we introduce the following technical innovations:
  • Hierarchical prototype learning strategy (C1). Traditional prototypes are typically limited to explaining individual classification results, while we extend the prototypes so that they model a temporal pattern, which consists of a sequence of regression results. Learning such a sequential temporal pattern is hard to balance predictive performance with interpretability. Using too few prototypes leads to poor forecasting accuracy, while employing too many can compromise interpretability by producing overly similar patterns. To address this, we design a hierarchical prototype learning strategy, where a small set of coarse prototypes ensures global interpretability, and prototypes are progressively refined into lower levels to capture local details and improve accuracy. This design allows experts to intervene by specifying which prototypes should be further split, enabling efficient and user-friendly model adjustments.
  • Multi-channel prototype similarity computing (C2). To enable effective modeling of the interactions among input variables of different types and their impact on the output curve patterns, we compute the prototype-instance similarity by incorporating a multi-channel embedding and bottleneck fusion mechanism, which improves both the interpretability and the accuracy of the model.
Our framework achieves state-of-the-art performance with good interpretability and steerability. In particular, ProtoTS reduces MSE by 48.3% and MAE by 20.9% compared to the state-of-the-art model on the LOF dataset. Moreover, its accuracy decreases much more slowly when the amount of training data is reduced, compared to the baselines. A case study demonstrates that our explanations enable an easy-to-understand overall and detailed understanding. Expert edits based on the explanations reduced MSE by 0.009 and further improved explainability.
2 Related Work
Time Series Forecasting with Exogenous Variables Time series forecasting with exogenous variables has been extensively explored in both classical and modern approaches. Traditional statistical models such as ARIMAX (williams2001multivariate) and SARIMAX (vagropoulos2016comparison) extend the ARIMA framework to incorporate correlations between exogenous and endogenous variables. Seeking richer dynamics, Transformer families like TFT (lim2021temporal) and TimeXer (wang2024timexer), introduce attention mechanisms to tackle this task. Parallel work in MLPs pursues efficiency and long-horizon stability: NBEATSx (oreshkin2019n) appends a dedicated stack for auxiliary factors on top of a residual basis expansion, TiDE (das2023long) factorizes temporal and feature projections with two dense encoders to harness known‑future variables, and TSMixer (ekambaram2023tsmixer; chen2023tsmixer) alternates mixing layers to effectively model their interactions. Outside deep learning, generic tabular learners such as XGBoost (chen2016xgboost), LightGBM (ke2017lightgbm), and GAM (yang2023interactive), perform well when exogenous variables are informative and strongly correlated with the target. Our method, ProtoTS, achieves strong performance on this task while offering interpretability and steerability for domain experts, effectively handling large amounts of heterogeneous variables while filtering out irrelevant information.
Time Series Interpretability Time series interpretability can be divided into post-hoc and ante-hoc approaches. Post-hoc interpretability focuses on unveiling the reasoning of pre-existing black-box models (gu2025explainable). Generic explanation techniques such as SHAP (lundberg2017unified), LIME (ribeiro2016should), IG (sundararajan2017axiomatic) can be directly applied to time-series models. crabbe2021explaining learns sparse perturbation masks over the time-feature dimensions, struggling to quantify each input’s effect on a regression forecast. Consequently, regression tasks tend to favour ante-hoc models whose interpretability is built in by design. Following the attribution-centric view of post‑hoc explanations, TFT (lim2021temporal) provides explorable attention over input time steps, and DiPE-Linear (zhao2024disentangled) parameterizes filters in both temporal and spectral domains to visualise the regions of interest. N-BEATS (oreshkin2019n) uses residual blocks for generic, trend, and seasonal components, and olivares2023neural extends the scheme to exogenous variables. These works are limited to local explanations and fail to provide a global view that experts need for a shared understanding. CycleNet (lin2024cyclenet) identifies common periodic patterns in time-series data but only focuses on endogenous variables. In contrast, ProtoTS holds a global set of prototypical patterns and explains each forecast by matching it to them, offering both global overview and local interpretation. Its hierarchical structure supports expert steering at multiple levels, ensuring faithful overall understanding.
Prototypical Time Series Model Previous works have already shown the value of prototypes in time series models. These methods typically fall into two categories. 1) prototypes decoded as intermediate representations (shen2025learn; li2023prototype; jin2023time) or model parameters (chen2024similarity). For example, jin2023time uses prototypes to map time-series embeddings into textual embeddings for large language models. Such designs enrich the model’s architecture but leave the prototypes detached from the output. 2) prototypes directly mapped to output variables (queen2023encoding; liu2024timex++; ming2019interpretable; obermair2023example), serving as typical examples of the outputs. While the latter offers stronger interpretability, existing approaches decode a prototype into only a single output variable (e.g., a class prediction). In contrast, our method is, to the best of our knowledge, the first to decode a prototype into a sequence of output variables (e.g., a 96-step forecasting curve), enabling holistic interpretation of temporal dynamics.
3 Method
Figure 2: The overall framework of ProtoTS, which comprises two main modules: the multi-channel prototype similarity computation module and the hierarchical prototype learning module.
3.1 Formulation and Framework
Problem Formulation We consider the task of time series forecasting with exogenous variables, as shown in Figure 1(a), since abundant heterogeneous exogenous variables provide rich information to boost endogenous predictability. Formally, let the historical endogenous variables (e.g., electric load) as $\mathbf{Y}_{1:L}=\{\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{L}\}\in\mathbb{R}^{L\times 1}$ and the associated exogenous variables (e.g., temperature and holiday) as $\mathbf{X}_{1:L}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{L}\}\in\mathbb{R}^{L\times C}$, where $L$ is the look-back window size and $C$ is the number of exogenous variables. In many practical scenarios (wessel2020using; lin2023prediction), future exogenous variables $\mathbf{X}_{L+1:L+H}=\{\mathbf{x}_{L+1},\dots,\mathbf{x}_{L+H}\}\in\mathbb{R}^{H\times C}$ are also available. For example, in the LOF dataset, we obtain forecasted weather variables (e.g., temperature) from commercial weather services, and calendar variables (e.g., holiday flags) are accessible in advance. In our setting, we explicitly incorporate future exogenous variables, as these substantially improve forecasting performance. The goal is to predict the future curve of the endogenous variable for the next $H$ time steps, denoted as $\mathbf{Y}_{L+1:L+H}=\{\mathbf{y}_{L+1},\dots,\mathbf{y}_{L+H}\}\in\mathbb{R}^{H\times 1}$. The forecasting model $\mathcal{F}$ takes the input variables to predict the future endogenous variables $\widehat{\mathbf{Y}}$:
$\widehat{\mathbf{Y}}_{L+1:L+H}=\mathcal{F}\Big(\mathbf{Y}_{1:L},\mathbf{X}_{1:L},\mathbf{X}_{L+1:L+H}\Big).$ (1)
Framework Overview The ProtoTS framework is illustrated in Figure 2. It contains two modules. The multi-channel prototype similarity computation module effectively models the interactions between input variables and their impact on prototypes. The hierarchical prototype learning module structures prototypes hierarchically, with coarse-grained root prototypes capturing global patterns and fine-grained child prototypes modeling local variations, ensuring good explainability, steerability, and accuracy. The two modules will be explained in detail in Sections 3.2 and 3.3.
3.2 Multi-Channel Prototype Similarity Computation
This module ensures effective modeling of interactions between input variables and their relations with prototypes by using three steps: the multi-channel embedding step effectively embeds heterogeneous variables in separate channels, the bottleneck channel fusion step then models complex interactions between variables without causing overfitting and outputs the final input embedding, and the prototype similarity computing step finally estimates the relations between each input and the prototypes by computing the similarities between their embeddings.
Multi-Channel Embedding In ProtoTS, we adopt a multi-channel embedding where endogenous and exogenous variables are processed separately to achieve optimal encoding for each type. Specifically, endogenous variables are directly encoded through an endogenous encoding channel, implemented as a non-linear projection $\gamma$ using a multi-layer perceptron with activation functions, a widely used design (das2023long). For exogenous variables, we follow standard practice in tabular learning (wang2022transtab). We categorize them into discrete variables (e.g., is holiday, day of week) and continuous variables (e.g., temperature, humidity), and apply tailored embedding methods to best capture the properties of each variable. Concretely, exogenous variables $\mathbf{x}_{t}$ at time $t$ are decomposed into $\mathbf{x}_{t}^{\text{dis}}\in\mathbb{N}^{C_{\text{dis}}}$ and $\mathbf{x}^{\text{con}}_{t}\in\mathbb{R}^{C_{\text{con}}}$, $C_{\text{dis}}$ and $C_{\text{con}}$ refer to the number of discrete and continuous variables. Discrete variable $\mathbf{x}_{t,j}^{\text{dis}}$ are processed through a discrete encoding channel using dedicated embedding tables $\mathbf{E}_{j}\in\mathbb{R}^{|\Omega_{j}|\times d}$, where $\Omega_{j}$ is the vocabulary and $d$ is the embedding dimension. Continuous variables $\mathbf{x}^{\text{con}}_{t,j}$ are projected into the embedding space through a continuous encoding channel, applying variable-specific non-linear projections $\psi_{j}$. This multi-channel embedding allows the model to better preserve the information contained in heterogeneous variables.
Bottleneck Channel Fusion To effectively fuse all variables information at time step $t$ and obtain a full representation, we aggregate the embeddings of all variables through addition, which integrates information across heterogeneous variables without additional parameters (arora2017simple):
$\mathbf{Z}_{t}=\begin{cases}\gamma(\mathbf{y}_{t})+\sum\limits_{j=1}^{C_{\text{dis}}}\mathbf{E}_{j}(\mathbf{x}^{\text{dis}}_{t,j})+\sum\limits_{j=1}^{C_{\text{con}}}\psi_{j}(\mathbf{x}^{\text{con}}_{t,j}),&t\in[1:L]\quad\text{ (Look-back Window)}\\ \sum\limits_{j=1}^{C_{\text{dis}}}\mathbf{E}_{j}(\mathbf{x}^{\text{dis}}_{t,j})+\sum\limits_{j=1}^{C_{\text{con}}}\psi_{j}(\mathbf{x}^{\text{con}}_{t,j}),&t\in[L+1:L+H]\quad\text{ (Forecast Window)}\end{cases}$ (2)
During the forecast window, where $\mathbf{y}_{t}$ is the target to be predicted, only exogenous variables are used to construct $\mathbf{Z}_{t}$. The embedding $\mathbf{Z}_{1:L+H}=\{\mathbf{Z}_{1},\cdots,\mathbf{Z}_{t},\cdots,\mathbf{Z}_{L+H}\}\in\mathbb{R}^{(L+H)\times d}$ now encodes a rich amount of information from endogenous and exogenous variables. However, such a rich embedding may also introduce noise, as irrelevant information from heterogeneous inputs can be mixed into the representations. To mitigate this, ProtoTS introduces a bottleneck layer to filter out irrelevant information while retaining the most predictive components. Concretely, we project the high-dimensional embeddings into a lower-dimensional latent space through a bottleneck projection, $\mathbb{R}^{d}\to\mathbb{R}^{d_{\text{bottle}}}\to\mathbb{R}^{d}$, where $d_{\text{bottle}}\ll d$. The compressed representation is then mapped back to the original space, retaining the most relevant information while filtering out noise (wang2021variational). We apply the bottleneck layer within an MLP-Mixer architecture (tolstikhin2021mlp), inserting it into both $\text{MLP}_{\text{feature}}$ and $\text{MLP}_{\text{time}}$. Fusion is performed sequentially, first along the feature dimension and then along the temporal dimension, T denotes the transpose of the vector:
$\mathbf{Z}_{1:L+H}^{(l+1)}=\text{MLP}_{\text{time}}(\text{MLP}_{\text{feature}}(\mathbf{Z}_{1:L+H}^{(l)})^{\text{T}})^{\text{T}}.$ (3)
The bottleneck fusion layer is stackable, where $l$ denotes the layer index. After stacking, we obtain $\mathbf{Z}_{1:L+H}\in\mathbb{R}^{(L+H)\times d}$ in the original embedding dimension. To enable prototype matching in a unified one-dimensional vector space, we perform a linear aggregation along the temporal dimensions:
$\hat{\mathbf{Z}}=\mathbf{Z}_{1:L+H}^{\text{T}}\text{W},\hat{\mathbf{Z}}\in\mathbb{R}^{d}$ (4)
where $\text{W}\in\mathbb{R}^{(L+H)\times 1}$ projects the sequence length from $L+H$ to 1.
Prototype Similarity Computing To provide a user-friendly global explanation while handling complex information at a local level for detailed interpretation and accurate forecasting, ProtoTS integrates prototypical explanation directly into its prediction mechanism via a learnable set of prototypes $\mathbf{\Pi}$. Each prototype in $\mathbf{\Pi}$ is represented as an embedding $\bm{\mu}\in\mathbb{R}^{d}$ and its corresponding temporal pattern $\mathbf{p}\in\mathbb{R}^{T}$, which are both learnable parameters. Here, $\bm{\mu}$ is used to compute the similarity between the query representation $\hat{\mathbf{Z}}$ and the prototypes, by applying a softmax over their distances $d(\hat{\mathbf{Z}},\bm{\mu})$. The distance can be chosen flexibly, such as Euclidean distance, cosine similarity. We adopt Euclidean distance, under which the prototypes are equivalent to a Gaussian mixture model in the embedding with an identity covariance matrix (allen2019infinite). Meanwhile, temporal pattern $\mathbf{p}$ is used to form the final prediction. Specifically, we fix each prototype temporal pattern $\mathbf{p}$ over a predefined period of length $T$, and during forecasting, the prototype pattern is aligned with the forecast window $H$ based on its phase and length (lin2024cyclenet).
3.3 Hierarchical Prototype Learning
The key challenge in designing the prototype set $\mathbf{\Pi}$ lies in balancing predictive performance with interpretability. Using too few prototypes may lead to poor forecasting accuracy, while employing too many can compromise interpretability by introducing overly similar patterns that are hard to distinguish and reason about. To address this trade-off, ProtoTS adopts a hierarchical prototype learning strategy that organizes the prototype set in a tree structure, effectively combining modeling capacity with multi-level interpretability.
Root Level Learning At the root level, ProtoTS aims to capture the overall pattern using only a small set of prototypes, thus ensuring interpretability by allowing users to efficiently form a global understanding. These root prototypes typically correspond to coarse-grained temporal patterns that consistently recur over long horizons (e.g., seasonal trends, holiday effects; see Figure 4(Layer-1)). The similarity of root prototype $c$ is computed based on $\bm{\mu}_{c}$, followed by a softmax operation:
$f(\hat{\mathbf{Z}}|\bm{\mu}_{c})=\frac{\exp(-d(\hat{\mathbf{Z}},\bm{\mu}_{c}))}{\sum_{i=1}^{N}\exp(-d(\hat{\mathbf{Z}},\bm{\mu}_{i}))}.$ (5)
The final prediction $\widehat{\mathbf{Y}}$ is then generated as a weighted combination of the temporal patterns from all root prototypes:
$\widehat{\mathbf{Y}}_{L+1:L+H}=\sum_{i=1}^{N}f(\hat{\mathbf{Z}}|\bm{\mu}_{i})\cdot\mathbf{p}_{i}.$ (6)
At the first stage, the model is trained using the root prototypes until convergence. This root-level learning allows the model to establish a globally interpretable set of representative temporal patterns, providing users with an intuitive understanding of the model’s decision logic at a high level.
Splitting Strategy Once the model converges, each root prototype is further split into $M$ child prototypes to introduce fine-grained variations. For root prototype $i$, its $M$ child prototypes are denoted as $\{(\bm{\mu}_{i,j},\mathbf{p}_{i,j})\}_{j=1}^{M}$. The similarity between $\hat{\mathbf{Z}}$ and its child prototypes is computed within this local group, ensuring all child prototypes remain associated with their root:
$f(\hat{\mathbf{Z}}|\bm{\mu}_{i,k})=\frac{\exp(-d(\hat{\mathbf{Z}},\bm{\mu}_{i,k}))}{\sum_{j=1}^{M}\exp(-d(\hat{\mathbf{Z}},\bm{\mu}_{i,j}))}.$ (7)
Based on this hierarchical matching, the final prediction is formed as a weighted combination of all child prototypes, modulated by both root-level and child-level similarities. The complete forecasting output is given by:
$\widehat{\mathbf{Y}}_{L+1:L+H}=\sum_{i=1}^{N}f(\hat{\mathbf{Z}}|\bm{\mu}_{i})\sum_{j=1}^{M}f(\hat{\mathbf{Z}}|\bm{\mu}_{i,j})\cdot\mathbf{p}_{i,j}.$ (8)
While the above describes a two-level hierarchy, the structure is general: any leaf prototype can be incrementally split to form a multi-level tree. However, not all prototypes need splitting, and some splits may yield overly similar temporal patterns. Therfore, we propose a splitting rule for determining which leaf prototypes should be further refined. The rule is designed to identify prototypes whose current temporal patterns are insufficiently representative of their associated instances and thus require additional fine-grained refinement. The pseudocode is provided in Algorithm 1.
Algorithm 1 Prototype Splitting Rule
Input: Training dataset $\mathcal{D}$, leaf prototypes $\Pi_{\text{leaf}}=\{\text{leaf}_{1},\dots\text{leaf}_{n_{\text{leaf}}}\}$, top-$k$ activation count, selection ratio $\alpha$.
Initialize: $\text{Loss}[l]\leftarrow 0$, $\text{Count}[l]\leftarrow 0$ for all leaf prototypes $\text{leaf}_{l}$.
For each training instance $(x,y)$ in $\mathcal{D}$:
[…]
Compute similarity scores $\{s_{l}\}_{l=1}^{n_{\text{leaf}}}$ over all leaf prototypes.
Let $\mathcal{T}$ be indices of top-$k$ similar prototypes based on $s_{l}$.
For each $l\in\mathcal{T}$:
$\text{Loss}[l]\leftarrow\text{Loss}[l]+L$.
$\text{Count}[l]\leftarrow\text{Count}[l]+1$.
Normalized loss:
$\text{NormLoss}[l]=\begin{cases}\frac{\text{Loss}[l]}{\text{Count}[l]},&\text{if }\text{Count}[l]>0,\\ 0,&\text{otherwise}.\end{cases}$
Select leaf prototypes with top $\alpha\%$ values in $\text{NormLoss}[l]$.
This hierarchical design enables experts to first obtain a global understanding through high-level root prototypes (e.g., distinguishing between seasonal or holiday-related patterns), and then refine their insights by examining lower-level child prototypes that capture more localized variations (e.g., differentiating between long holidays and short holidays). Moreover, the hierarchy empowers expert interaction. Experts can steer the model by selectively splitting specific prototypes, introducing new root-level prototypes, or directly editing the temporal pattern to integrate domain knowledge.
Loss Function In addition to the standard L1 forecasting loss, we incorporate an entropy-based regularization on the prototype weights $f(\hat{\mathbf{Z}}\mid\bm{\mu}_{i})$, with $\lambda$ controlling the strength, to encourage a few main prototypes to cover most predictions. We note that an L2 loss can also be used.
[…]
4 Experiments
We conduct extensive experiments to thoroughly evaluate the effectiveness of ProtoTS. The main results show that ProtoTS consistently delivers strong performance on time series forecasting with exogenous variables. ProtoTS also achieves high accuracy on the multivariate time series forecasting, as shown in Appendix G. We further analyze the effectiveness of model components and the quantitative evaluation confirms that ProtoTS provides the most understandable and usable explanation. Finally, a case study illustrating our ability to provide global explanations and fine-grained understanding, while supporting expert editing to enhance both performance and interpretability.
4.1 Overall Performance
Datasets Load Forecasting dataset (LOF) is a synthetic electric load dataset, including 22 supporting covariates and covering four regions. We also evaluated on the Electricity Price Forecasting (EPF) dataset (lago2021forecasting), which covers hourly electric prices from five Nordic markets. The specific details of the datasets can be found in the Appendix B.1.
Baselines To provide a wide‑ranging comparison, we benchmark eight models in total: five state-of-the-art deep-learning methods (TimeXer (wang2024timexer), iTransformer (liu2023itransformer), TiDE (das2023long), TFT (lim2021temporal), N‑BEATSx (olivares2023neural)) together with three strong machine-learning baselines (XGBoost (chen2016xgboost), LightGBM (ke2017lightgbm), pyGAM (serven2018pygam)). All deep-learning models support exogenous variables as inputs, and we further extend TimeXer and iTransformer to incorporate future covariates. The machine learning models generate forecasts recursively with exogenous inputs.
Results The MAE on the LOF dataset is shown in Table 1, ProtoTS achieves state-of-the-art performance across four datasets, reducing MSE by 48.3% and MAE by 20.9% over the best baseline. Interpretable models like N-BEATSx are sensitive to outliers, resulting in high MSE. Transformer-based models fail to capture heterogeneous exogenous information due to single-channel encoding. Although TiDE incorporates exogenous variables in both encoder and decoder, its performance is hindered by irrelevant noise. Machine learning models perform well on stable datasets but struggle with complex scenarios. In contrast, ProtoTS effectively captures diverse information and filters noise, achieving superior performance. Visual comparison can be found in Appendix E.
Table 1: MAE results on the LOF dataset with 22 supporting covariates. The look-back window and forecast window are set to 384 and 96 for all methods. $\Delta$ means the relative improvement of ProtoTS over other baselines. Full MSE results can be found in Appendix C.1.
Model  ProtoTS TimeXer  iTrans.   TiDE   TFT NBEATSx XGBoost LightGBM pyGAM
LOF RE 0.198 0.272 0.279 0.253 0.342 0.388 0.405 0.366 0.388
YC 0.055 0.079 0.080 0.057 0.116 0.572 0.084 0.082 0.117
EA 0.059 0.096 0.097 0.061 0.108 0.589 0.092 0.085 0.108
PC 0.112 0.182 0.139 0.164 0.285 0.704 0.230 0.222 0.269
Avg 0.106 0.157 0.149 0.134 0.213 0.563 0.203 0.189 0.221
$\Delta$ - 32% 29% 21% 50% 81% 48% 44% 52%
We also evaluate ProtoTS on the EPF dataset. Unlike LOF, this task involves fewer covariates but exhibits higher volatility, posing challenges for accurate forecasting. As shown in Table 2, ProtoTS achieves the best performance across all five markets, reducing MSE by 8% and MAE by 8% compared to the best-performing baseline. We further conduct a sensitivity analysis under 5 random seeds, see Appendix C.3. Full implementation details can be found in the Appendix B.2.
Table 2: MAE results on EPF dataset equipped with 6 supporting covariates. The look-back window and forecast window are set to 168 and 24 for all methods. $\Delta$ means the relative improvement of ProtoTS over other baselines. Full MSE results can be found in Appendix C.2
Model  ProtoTS TimeXer  iTrans.   TiDE   TFT NBEATSx XGBoost LightGBM pyGAM
EPF NP 0.213 0.240 0.264 0.318 0.277 0.266 0.488 0.431 0.480
PJM 0.152 0.173 0.171 0.226 0.226 0.182 0.292 0.256 0.351
BE 0.226 0.241 0.266 0.323 0.277 0.324 0.451 0.392 0.607
FR 0.183 0.192 0.213 0.281 0.242 0.340 0.374 0.361 0.570
DE 0.318 0.343 0.335 0.474 0.489 0.408 0.539 0.506 0.551
Avg 0.218 0.238 0.250 0.324 0.302 0.304 0.429 0.389 0.512
$\Delta$ - 8% 13% 33% 28% 28% 49% 44% 57%
4.2 Ablation Study and Sensitivity Analysis
Ablation Study We conduct an ablation study to assess the contribution of the three key components in ProtoTS: multi-channel embedding, bottleneck layer and hierarchical structure. To assess the modeling of heterogeneous variables, we perform ablations by simplifying the embedding to a single channel, removing the bottleneck layer, and flattening the hierarchical prototype structure. The ablation results in Table 3 demonstrate the importance of each key component in ProtoTS. While removing any component degrades performance, the full model with all three achieves the best results, confirming the effectiveness of the overall design.
Table 3: Ablation Results. Evaluating the performance after removing one of the following components: bottleneck layer, multi-channel embedding, and hierarchical structure.
Models PC YC RE EA AVG
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
w/o bottleneck 0.044 0.160 0.013 0.089 0.129 0.248 0.010 0.073 0.049 0.143
w/o multi-channel 0.034 0.130 0.006 0.058 0.108 0.216 0.007 0.062 0.039 0.117
w/o hierachy 0.026 0.117 0.006 0.060 0.089 0.202 0.007 0.061 0.032 0.110
ProtoTS 0.025 0.112 0.006 0.055 0.085 0.198 0.007 0.059 0.031 0.106
Number of Root Prototypes We evaluate how the number of root prototypes affects model performance, as shown in Figure 3(a). The results clearly show that prototypes number is correlated with forecasting accuracy. Increasing the number of prototypes leads to lower MAE, as more prototypes enable the model to capture a richer variety of temporal patterns for better representation. However, once the number of root prototypes reaches a threshold, the performance no longer improves. This indicates that the prototypes have already covered all typical patterns in the dataset.
Proportion of Train Data To evaluate the impact of limited training data on model performance, we gradually increase the size of the PC dataset from 50% to 100%. As the available data decreased, existing baselines such as TimeXer, iTransformer, and TiDE show clear drops in MSE. In contrast, ProtoTS maintains stable results with only minor performance loss, which improves high data efficiency. The comparison results are illustrated in Figure 3(b).
Figure 3: Sensitivity analysis: (a) Effect of increasing root prototype numbers from {6, 9, 12, 15, 18}. (b) Data efficiency as the training data proportion increases from 50% to 100%.
Table 4: Quantitative Evaluation.
Model UPrec $\uparrow$ SUS Score $\uparrow$
ProtoTS 77 $\pm$ 3.6% 73.36
TFT 64 $\pm$ 3.4% 29.74
NBEATSx 62 $\pm$ 2.8% 38.66
4.3 Quantitative Evaluation of Interpretability
In this section, we conduct a quantitative evaluation of interpretability. We aim to answer the two questions: 1) How understandable and accurate are the prototypes in explaining the predictions? 2) How usable are the explanation systems provided by different interpretable models (ProtoTS, TFT, NBEATSx)? Specifically, we provide 24 users with different explanations as well as the corresponding predictions, where the three models are presented randomly to avoid bias. Users were asked concrete questions about the input variables, such as "Given this prediction and explanations, which season does the predicted day most likely belong to?". To evaluate interpretability, we define User Precision (UPrec) as the proportion of correctly answered variable-related questions, reflecting whether users can accurately infer the values of relevant input features from the explanations. To further assess usability, we adopt the System Usability Scale (SUS) (brooke1996sus), a widely used questionnaire that measures users’ perceived ease of use and overall satisfaction with a system. The quantitative results are reported in Table 4, with additional details in Appendix D.
4.4 Case Study of Interpretability and Steerability
Figure 4 illustrates how ProtoTS discovers meaningful and interpretable hierarchical prototypes that help better understand electricity load forecasting and facilitate expert edits that improves model performance. We use the RE case in LOF. The load is shaped by multiple external factors, including seasonal, temperature and weekday/holiday. ProtoTS is initialized with 6 root prototypes, each further refined into 2 child prototypes. This hierarchy captures both coarse and fine temporal patterns. Time series predictions are weighted combinations of these prototypes (Figure 4(a)). The activation weights link each prototype’s contribution to predictions and are also the key to interpretation.
Figure 4: Case study of interpretability and steerability: (a) Activation patterns of prototypes, which each colored bar shows the degree of activation of its corresponding prototype each day. This visualization shows seasonal, weekly, and holiday-related prototype activations. (b) The learned prototype hierarchy that illustrates typical temporal patterns and their correlations with covariates for interpretable load forecasting. After the expert splits the first prototype at layer-2, the mean MSE across five random seeds decreases from 0.099 to 0.090 for Spring Festival related predictions.
Obtaining overall understanding of temporal patterns. The root prototypes provide coarse temporal patterns of high-level concepts (Figure 4(b), layer-1): workdays, non-workdays, holidays, winter, mild temperature, and summer. For example, the summer pattern (blue) exhibits high daytime peaks due to air-conditioning dominance. Autumn and winter data are dominated by distinct seasonal load profiles in red and orange. The brown prototype with a dual-peak pattern is active during holidays. This contrasts with the green weekday pattern with a single afternoon peak. The dual peaks suggest that industrial activity is paused, and the load is driven by residential behavior.
Fine-grained analysis of detailed temporal patterns. The child prototypes provide fine-grained refinements of these general patterns (Figure 4(b), layer-2). For example, the root holiday prototype is split into two children: one capturing the pattern of the major holiday, Spring Festival (Chinese New Year), and the other representing the rest of the normal holidays. The Spring Festival prototype shows a more pronounced evening peak and flattened noon peak.
Steering the model to improve accuracy. Editing the model for improvements becomes feasible thanks to the clearly interpretable temporal patterns learned in protoTS (Figure 4(b), layer-3). Based on domain knowledge that the Spring Festival itself contains distinct sub-patterns of pre- and mid-holiday behaviors, we manually split the Spring Festival prototype into two prototypes, which learns a base pattern and a special pattern. The base pattern stays consistently active through the whole Spring Festival period. The special pattern starts from the Lunar Jan. $1^{\text{st}}$, strengthening progressively day by day, reflecting a deepening holiday atmosphere: reduced daytime electrical usage, and stronger evening peaks as people stay home. This edit reduced the MSE during Spring Festival by 0.009, showing how expert adjustments can enhance both accuracy and explainability.
5 Conclusion
In this work, we proposed ProtoTS, a novel forecasting framework that achieves both high accuracy and multi-level interpretability through hierarchical prototypes. It learns low-noise representations from heterogeneous variables and matches them with prototypes to generate predictions and explanations, while the hierarchy enables coarse-to-fine pattern learning and expert-steerability. Extensive experiments and case studies on realistic datasets show state-of-the-art performance and transparent support, enabling experts to refine model behavior and bridging accuracy with interpretability.
Ethics statement
We provide an ethics statement as follows:
  • Human subjects research: Our experiments include a user study (see Section 4.3). The study does not raise any ethical concerns, and all participants gave informed consent prior to participation.
  • Datasets: All datasets used in our work are publicly available. The released dataset is a simulation of real-world data and does not contain any actual sensitive information.
  • Fairness and bias: Our work does not raise fairness or bias concerns.
Reproducibility statement
We confirm that the contents of the main text and the appendix are sufficient to reproduce our work and our experiments are reproducible. Section 3 clearly describes the proposed method and workflow. Appendix B.1 provides detailed dataset descriptions, and Appendix B.2 provides sufficient implementation details (including hyperparameters and training settings). We also include in the abstract a link to an anonymous repository containing our source code and the released dataset; the repository provides scripts for data preparation, training, and evaluation, along with configuration files and random seeds. We have verified that the code is consistent with the descriptions in the paper and that the reported experiments can be reproduced following the provided instructions.
Appendix A Use of LLMs
We used large language models (LLMs) only to aid or polish writing. Specifically, LLMs were employed to: (1) check grammar and spelling; (2) assist with professional academic English phrasing without hurting the original meaning; and (3) perform light-length reduction to meet formatting constraints. LLMs were not used to generate ideas, methods, experiments, figures, or results. LLMs were not used to find related works or to retrieve references. All suggestions were reviewed and verified by the authors, who take full responsibility for the final text.
Appendix B Implementation Details
B.1 Dataset Descriptions
LOF dataset The LOF dataset is an synthetic electric load forecasting dataset. It is designed to predict future electricity demand at the provincial and city levels, enable electricity generators, distributors, and suppliers to plan effectively ahead and promote energy conservation among the users (nti2020electricity). Due to the influence of human industrial and commercial activities, electricity demand exhibits a clear daily cycle, maintaining a relatively stable and predictable pattern under normal conditions. The difficulty of load forecasting mainly arises from special instances that are strongly correlated with a large amount of exogenous information, such as time-related attributes and comprehensive weather data. Consequently, accurate electricity load forecasting requires models to not only effectively handle these variables but also provide insights into how they affect prediction results. Hence, electricity load forecasting is a typical task of time series forecasting with exogenous variables, characterized by a high demand for interoperability. The LOF dataset ranges across three years, with a sampling interval of 15 minutes. The LOF dataset includes 22 exogenous variables, consisting of 14 discrete time-related variables and 8 continuous weather variables. The specific details of these exogenous variables are presented in Table 10.
EPF dataset The EPF dataset is an electricity price forecasting datasets (lago2021forecasting). It collects from five distinct day-ahead electricity markets in Northern Europe, covering a six year period with a sampling interval of 1 hour. The five datasets are described as follows: (1) NP corresponds to the Nord Pool market, providing hourly electricity prices, grid load data, and wind power forecasts from 2013-01-01 to 2018-10-24. (2) PJM refers to the Pennsylvania-New Jersey-Maryland market, including zonal electricity prices for the Commonwealth Edison (COMED) zone, along with system-wide load and COMED-specific load forecasts, covering the period from 2013-01-01 to 2018-10-24. (3) BE represents the Belgian electricity market, offering hourly price data, load forecasts for Belgium, and generation forecasts for France, collected from 2011-01-09 to 2016-10-31. (4) FR pertains to the French electricity market, comprising hourly price records, load forecasts, and generation forecasts over the timeframe from 2012-01-09 to 2017-10-31. (5) DE covers the German electricity market, documenting hourly electricity prices, zonal load forecasts for the TSO Amprion area, as well as wind and solar generation forecasts, with data available from 2012-01-09 to 2017-10-31.
B.2 Implementation Details
All experiments are implemented using PyTorch (paszke2019pytorch) and executed on a single NVIDIA V100 GPU with 32GB of memory. For model optimization, we employ the Adam (kingma2014adam) optimizer with an initial learning rate of and use our Loss as the objective function. The training is set for a maximum of 30 epochs, with early stopping applied to prevent overfitting. In our proposed model, the number of bottleneck fusion blocks is selected from the set {1, 2, 3}, while the dimension of the series representations, denoted as model dimensions, is chosen from {32, 64, 128, 256, 512}, the number of initial prototypes is searched from {6, 8, 12}, the deepest level is chosen from {1, 2, 3}, the number of each prototype can split is selected from the set {2, 3}. All baseline models used for comparison are re-implemented based on the TimeXer Repository benchmark. All baseline use REVIN (kim2021reversible) in training and test process.
Appendix C Full Results
C.1 Full Results of LOF dataset
Full results of the LOF dataset can be found in Table 5, including both MSE and MAE metrics. ProtoTS achieves state-of-the-art performance across all four datasets, reducing MSE by 48.3% and MAE by 20.9%. TimeXer, iTransformer, and TiDE show the most competitive results, with TiDE performing the best among them. Machine learning methods achieve decent performance with very short training time.
Table 5: Full results of LOF dataset on the time series forecasting with exogenous variables task, equipped with 22 supporting covariates. The look-back window and forecast window are set to 384 and 96 for all baselines. $\Delta$ means the relative improvement of ProtoTS over other baselines.
Model RE YC EA PC AVG $\Delta$
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ProtoTS 0.085 0.198 0.006 0.055 0.007 0.059 0.025 0.112 0.031 0.106 - -
TimeXer 0.167 0.272 0.011 0.079 0.015 0.096 0.073 0.182 0.067 0.157 53% 32%
iTransformer 0.169 0.279 0.013 0.080 0.018 0.097 0.043 0.139 0.061 0.149 49% 29%
TiDE 0.158 0.253 0.007 0.057 0.008 0.061 0.066 0.164 0.060 0.134 48% 21%
TFT 0.228 0.342 0.025 0.116 0.021 0.108 0.154 0.285 0.107 0.213 71% 50%
NBEATSx 0.815 0.388 1.553 0.572 2.185 0.589 2.467 0.704 1.755 0.563 98% 81%
XGBoost 0.300 0.405 0.014 0.084 0.017 0.092 0.101 0.230 0.108 0.203 71% 48%
LightGBM 0.253 0.366 0.012 0.082 0.014 0.085 0.094 0.222 0.093 0.189 67% 44%
pyGAM 0.263 0.388 0.022 0.117 0.019 0.108 0.126 0.269 0.108 0.221 71% 52%
C.2 Full Results of EPF dataset
Full results of the EPF dataset can be found in Table 6, including both MSE and MAE metrics. ProtoTS achieves state-of-the-art performance across all five datasets, reducing MSE by 8% and MAE by 8%. Similarly, TimeXer, iTransformer, and TiDE show competitive results, with TimeXer performing the best in this case. Machine learning methods perform poorly on EPF, where the number of covariates is limited.
Table 6: Full results of EPF dataset on the time series forecasting with exogenous variables task, equipped with 6 supporting covariates. The look-back window and forecast window are set to 168 and 24 for all baselines. $\Delta$ means the relative improvement of ProtoTS over other baseline.
Model NP PJM BE FR DE AVG $\Delta$
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ProtoTS 0.168 0.213 0.064 0.152 0.353 0.226 0.351 0.183 0.271 0.318 0.241 0.218 - -
TimeXer 0.194 0.240 0.078 0.173 0.365 0.241 0.355 0.192 0.319 0.343 0.262 0.238 8% 8%
iTransformer 0.208 0.264 0.076 0.171 0.375 0.266 0.386 0.213 0.282 0.335 0.265 0.250 9% 13%
TiDE 0.306 0.318 0.121 0.226 0.495 0.323 0.492 0.281 0.533 0.474 0.389 0.324 38% 33%
TFT 0.259 0.277 0.138 0.226 0.412 0.277 0.452 0.242 0.709 0.489 0.394 0.302 39% 28%
NBEATSx 0.222 0.266 0.079 0.182 0.507 0.324 1.920 0.340 0.430 0.408 0.632 0.304 62% 28%
XGBoost 1.002 0.488 0.276 0.292 0.817 0.451 0.622 0.374 0.618 0.539 0.669 0.429 64 % 49%
LightGBM 0.678 0.431 0.234 0.256 0.598 0.392 0.603 0.361 0.549 0.506 0.532 0.389 55 % 44%
pyGAM 0.599 0.480 0.281 0.351 1.028 0.607 1.116 0.570 0.671 0.551 0.739 0.512 67 % 57%
C.3 Sensitive Analysis
To investigate the consistency of the model’s performance across different randomized scenarios, we selected five random seeds and reported the mean and standard deviation of the model’s performance. The experiments in Table 7 and Table 8 show that the model maintains stable performance without statistically significant variations in most cases.
Table 7: Sensitive Analysis on LOF dataset. We report our performance in $\mathrm{mean}$$\pm$ $\mathrm{std}$ format.
Model ProtoTS TimeXer iTransformer TiDE
Metric MSE MAE MSE MAE MSE MAE MSE MAE
[…]
Table 8: Sensitive Analysis on EPF dataset. We report our performance in $\mathrm{mean}$$\pm$ $\mathrm{std}$ format.
Model ProtoTS TimeXer iTransformer TiDE
Metric MSE MAE MSE MAE MSE MAE MSE MAE
[…]
Appendix D User Study Details
We conducted the user study using a questionnaire format, the specific layout of which is shown in Figure 5. The questionnaire consists of two parts: Variable-related Questionnaire and System Usability Scale (SUS) Questionnaire. The first part is Variable-related Questionnaire, aims to evaluate the accuracy and comprehensibility of explanations provided by different explainable models. Participants are asked to perform causal inference on variables—such as holidays, seasons, Chinese New Year, and other related variables—based on the model’s predictions and its generated explanations. This section has unique, objective correct answers, and we measure the readability of the explanations by calculating the accuracy of participants’ responses. The second part is System Usability Scale (SUS) Questionnaire, designed to assess the usability of the explanation systems offering different levels of explainability. Participants are asked to subjectively rate the system based on their experience completing the first part. The SUS yields a single number representing a composite measure of the overall usability of the system under study. Note that scores for individual items are not meaningful on their own. To calculate the SUS score, first sum the score contributions from each item. Each item’s contribution ranges from 0 to 4. For items 1, 3, 5, 7, and 9, the contribution is the scale position minus 1. For items 2, 4, 6, 8, and 10, the contribution is 5 minus the scale position. Multiply the total sum of these contributions by 2.5 to obtain the overall SUS score, which ranges from 0 to 100.
Appendix E Visual Comparison
To intuitively illustrate the differences between ProtoTS and existing methods in predictions, we visualize the prediction results. The models chosen for comparison are iTransformer and Timexer, and the visualization is presented in Figure 6. For the prediction on August $2^{nd}$, ProtoTS successfully captures the evening peak, whereas the other two models perform poorly. For the predictions on October $9^{th}$ and December $10^{th}$, although all three models capture the overall shape reasonably well, ProtoTS preserves significantly more fine-grained details. Regarding the prediction for February $1^{st}$, ProtoTS more accurately forecasts noon peak of the day.
Appendix F Discussion
Although ProtoTS achieves high forecasting accuracy while providing multi-level, expert-steerable interpretability, it also has some potential limitations:
  • Limited adaptability to non-periodic patterns: ProtoTS relies on prototypical representations to capture typical temporal patterns. While this works well for seasonal or periodic data, its effectiveness diminishes when dealing with highly irregular, nonrepetitive time series. In such cases, more prototypes are needed to cover diverse patterns, increasing model complexity and reducing interpretability.
  • Dependence on domain expertise for refinement: The interpretability of ProtoTS benefits from expert-guided prototype editing. However, this process requires sufficient domain knowledge and manual effort. In domains where expert resources are hard to access or knowledge is difficult to formalize, the usability and effectiveness of this feature may be limited.
  • Computational cost: ProtoTS introduces additional computational overhead compared to transformer-based or MLP-based baselines (e.g., iTrans and TiDE). Under the same hardware and training configuration (optimizer, learning rate, sequence length, and batch size), ProtoTS shows higher per-iteration training time due to its multi-channel embedding design and hierarchical prototype modules, both of which require extra forward computations.
ProtoTS TimeXer iTrans TiDE
ms/iter 40.8 19.5 12.5 11.1
Table 9: Training time per iteration (ms).
Appendix G Multivariate Time Series Forecasting
In this section, we evaluate the performance of ProtoTS on standard multivariate time series forecasting tasks.
Datasets We conduct comprehensive experiments on six benchmark datasets to evaluate the multivariate forecasting capability of ProtoTS. These datasets cover a variety of domains:
  • ETTh1 and ETTh2 (zhou2021informer): Hourly electricity load data from two different nodes in a power grid. They capture seasonal patterns and variable interactions across multiple features, serving as standard benchmarks for long-term forecasting.
  • ETTm1 and ETTm2 (zhou2021informer): Minute-level electricity load datasets focusing on finer temporal granularity. They emphasize short-term dynamics and present challenges in modeling high-frequency fluctuations.
  • Weather (zhou2021informer): A multivariate weather dataset containing temperature, humidity, wind speed, and other meteorological variables collected from multiple stations.
  • Exchange (lai2018modeling): A financial time series dataset with daily exchange rates of eight foreign currencies.
For all datasets, we follow standard experimental protocols, using a 96 look-back window and forecasting multiple future steps {96, 192, 336, 720} to evaluate both short-term accuracy and long-horizon stability.
Baseline We compare ProtoTS against a broad set of baselines, including: Transformer-based models (TimeXer (wang2024timexer), iTransformer (liu2023itransformer), PatchTST (nie2022time), TimesNet (wu2022timesnet), Crossformer (zhang2023crossformer), Stationary (liu2022non), Autoformer (wu2021autoformer)), MLP-based architectures (TiDE (das2023long), DLinear (zeng2023transformers), RLinear (li2023revisiting), SCINet (liu2022scinet)).
Results The results in Table 11 indicate that ProtoTS achieves leading performance across most multivariate forecasting tasks. On structured datasets like ETTh1 and ETTh2, ProtoTS consistently outperforms all baselines at both short and long horizons, confirming its strength in modeling global temporal patterns. For high-frequency datasets such as ETTm1 and ETTm2, ProtoTS shows clear improvements in short-term forecasts, but its advantage narrows at longer horizons (e.g., 720 steps), where methods like PatchTST and TimesNet achieve comparable results. On the Weather dataset, ProtoTS maintains strong performance, though models like Crossformer remain competitive. For the challenging Exchange dataset, ProtoTS delivers robust forecasts but faces slight performance gaps against simpler models like DLinear at certain horizons. Overall, ProtoTS shows stable and strong performance across different tasks.
Figure 5: User Study Questionnaire. Bottom: An overview of the questionnaire’s purpose, question formats, and response instructions. Left: The Variable-related Questionnaire, where users are asked to infer input variables based on the provided explanations. Each question has a single correct answer, used to evaluate the accuracy of the system’s explanations. Right: The System Usability Scale (SUS) Questionnaire, which captures users’ subjective assessments to measure the usability of the current system.
Figure 6: Result Visualization: The prediction results for four sampled days, consistent with our experimental setting: Look-back window is 4 days and Predict window is 1 day, a day consists of 96 time steps. The horizontal axis represents time, and the vertical axis represents the current electricity load. The electricity load values have been normalized, so they are unitless and may include negative values.
Table 10: Descriptions of the exogenous variables used in the LOF dataset. The table includes each variable’s data type, whether it is discrete or continuous, its physical unit (if applicable), and a brief description of its meaning and role in the dataset.
Name Datatype Type Unit Description
Solar year int Discrete - gregorian calendar year
Solar month int Discrete - gregorian calendar month
Solar day int Discrete - gregorian calendar day
Lunar year int Discrete - chinese lunar calendar year
Lunar month int Discrete - chinese lunar calendar month
Lunar day int Discrete - chinese lunar calendar day
Hour int Discrete - hour of day
Minute int Discrete - minute of hour
Skin temperature float Continuous °C the temperature of the surface of the Earth
Surface pressure float Continuous hPa the temperature of sea water near the surface
Surface sensible heat flux float Continuous Jm-2 the transfer of heat between the Earth’s surface and the atmosphere through the effects of turbulent air motion
Total cloud cover float Continuous % the proportion of a grid box covered by cloud
Surface net solar radiation float Continuous Jm-2 the amount of solar radiation that reaches a horizontal plane at the surface of the Earth minus the amount reflected by the Earth’s surface
Total precipitation 4h lead float Continuous m the accumulated liquid and frozen water that falls to the Earth’s surface
10 metre wind float Continuous ms-1 maximum 3 second wind at 10 metre height as defined
Relative humidity float Continuous % the water vapour pressure as a percentage of the value at which the air becomes saturated
Is special workday bool Discrete - special workday added to make up for time lost during an adjusted holiday.
Is special holiday bool Discrete - public holiday mandated by law
Is holiday bool Discrete - rest day, including weekends and public holidays
Is workday bool Discrete - work day
Day of week int Discrete - monday=0, …, sunday=6
Pricing float Continuous - real‑time electricity price strategy
Table 11: Full results of multivariate time series forecasting task, where the input window is set to 96, and the forecast window is set to {96, 192, 336, 720}. Results are citepd from TimeXer (wang2024timexer), otherwise reproduced. $1^{\mathrm{st}}$ denotes the first count.
Model ProtoTS TimeXer iTrans. RLinear PatchTST Cross. TiDE TimesNet DLinear SCINet Stationary Auto.
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Etth1
96 0.380 0.388 0.382 0.403 0.386 0.405 0.386 0.395 0.414 0.419 0.423 0.448 0.479 0.464 0.384 0.402 0.386 0.400 0.654 0.599 0.513 0.491 0.449 0.459
192 0.437 0.423 0.429 0.435 0.441 0.436 0.437 0.424 0.460 0.445 0.471 0.474 0.525 0.492 0.436 0.429 0.437 0.432 0.719 0.631 0.534 0.504 0.500 0.482
336 0.483 0.440 0.468 0.448 0.487 0.458 0.479 0.446 0.501 0.466 0.570 0.546 0.565 0.515 0.491 0.469 0.481 0.459 0.778 0.659 0.588 0.535 0.521 0.496
720 0.484 0.459 0.469 0.461 0.503 0.491 0.481 0.470 0.500 0.488 0.653 0.621 0.594 0.558 0.521 0.500 0.519 0.516 0.836 0.699 0.643 0.616 0.514 0.512
AVG 0.445 0.427 0.437 0.437 0.454 0.447 0.446 0.434 0.469 0.454 0.529 0.522 0.541 0.507 0.458 0.450 0.456 0.452 0.747 0.647 0.570 0.537 0.496 0.487
Etth2
96 0.284 0.337 0.286 0.338 0.297 0.349 0.288 0.338 0.302 0.348 0.745 0.584 0.400 0.440 0.340 0.374 0.333 0.387 0.707 0.621 0.476 0.458 0.346 0.388
192 0.370 0.387 0.363 0.389 0.380 0.400 0.374 0.390 0.388 0.400 0.877 0.656 0.528 0.509 0.402 0.414 0.477 0.476 0.860 0.689 0.512 0.493 0.456 0.452
336 0.408 0.421 0.414 0.423 0.428 0.432 0.415 0.426 0.426 0.433 1.043 0.731 0.643 0.571 0.452 0.452 0.594 0.541 1.000 0.744 0.552 0.551 0.482 0.486
720 0.405 0.431 0.408 0.432 0.427 0.445 0.420 0.440 0.431 0.446 1.104 0.763 0.874 0.679 0.462 0.468 0.831 0.657 1.249 0.838 0.562 0.560 0.515 0.511
AVG 0.366 0.394 0.367 0.396 0.383 0.407 0.374 0.398 0.387 0.407 0.942 0.684 0.611 0.550 0.414 0.427 0.559 0.515 0.954 0.723 0.526 0.516 0.450 0.459
Ettm1
96 0.324 0.347 0.318 0.356 0.334 0.368 0.355 0.376 0.329 0.367 0.404 0.426 0.364 0.387 0.338 0.375 0.345 0.372 0.418 0.438 0.386 0.398 0.505 0.475
192 0.375 0.369 0.362 0.383 0.387 0.391 0.391 0.392 0.367 0.385 0.450 0.451 0.398 0.404 0.374 0.387 0.380 0.389 0.426 0.441 0.459 0.444 0.553 0.496
336 0.402 0.389 0.395 0.407 0.426 0.420 0.424 0.415 0.399 0.410 0.532 0.515 0.428 0.425 0.410 0.411 0.413 0.413 0.445 0.459 0.495 0.464 0.621 0.537
720 0.470 0.426 0.452 0.441 0.491 0.459 0.487 0.450 0.454 0.439 0.666 0.589 0.487 0.461 0.478 0.450 0.474 0.453 0.595 0.550 0.585 0.516 0.671 0.561
AVG 0.392 0.387 0.382 0.397 0.407 0.410 0.414 0.407 0.387 0.400 0.513 0.496 0.419 0.419 0.400 0.406 0.403 0.407 0.485 0.481 0.481 0.456 0.588 0.517
Ettm2
96 0.176 0.254 0.171 0.256 0.180 0.264 0.182 0.265 0.175 0.259 0.287 0.366 0.207 0.305 0.187 0.267 0.193 0.292 0.286 0.377 0.192 0.274 0.255 0.339
192 0.237 0.295 0.237 0.299 0.250 0.309 0.246 0.304 0.241 0.302 0.414 0.492 0.290 0.364 0.249 0.309 0.284 0.362 0.399 0.445 0.280 0.339 0.281 0.340
336 0.295 0.332 0.296 0.338 0.311 0.348 0.307 0.342 0.305 0.343 0.597 0.542 0.377 0.422 0.321 0.351 0.369 0.427 0.637 0.591 0.334 0.361 0.339 0.372
720 0.394 0.390 0.392 0.394 0.412 0.407 0.407 0.398 0.402 0.400 1.730 1.042 0.558 0.524 0.408 0.403 0.554 0.522 0.960 0.735 0.417 0.413 0.433 0.432
AVG 0.275 0.312 0.274 0.322 0.288 0.332 0.286 0.327 0.281 0.326 0.757 0.610 0.358 0.404 0.291 0.333 0.350 0.401 0.571 0.537 0.306 0.347 0.327 0.371
weather
96 0.155 0.197 0.157 0.205 0.174 0.214 0.192 0.232 0.177 0.218 0.158 0.230 0.202 0.261 0.172 0.220 0.196 0.255 0.221 0.306 0.173 0.223 0.266 0.336
192 0.202 0.241 0.204 0.247 0.221 0.254 0.240 0.271 0.225 0.259 0.206 0.277 0.242 0.298 0.219 0.261 0.237 0.296 0.261 0.340 0.245 0.285 0.307 0.367
336 0.259 0.285 0.261 0.290 0.278 0.296 0.292 0.307 0.278 0.297 0.272 0.335 0.287 0.335 0.280 0.306 0.283 0.335 0.309 0.378 0.321 0.338 0.359 0.395
720 0.337 0.333 0.340 0.341 0.358 0.349 0.364 0.353 0.354 0.348 0.398 0.418 0.351 0.386 0.365 0.359 0.345 0.381 0.377 0.427 0.414 0.410 0.419 0.428
AVG 0.238 0.264 0.241 0.271 0.258 0.279 0.272 0.291 0.259 0.281 0.259 0.315 0.271 0.320 0.259 0.287 0.265 0.317 0.292 0.363 0.288 0.314 0.338 0.382
Exchange
96 0.085 0.205 0.090 0.209 0.086 0.206 0.093 0.217 0.088 0.205 0.256 0.367 0.094 0.218 0.107 0.234 0.088 0.218 0.267 0.396 0.111 0.237 0.197 0.323
192 0.173 0.299 0.190 0.309 0.177 0.299 0.184 0.307 0.176 0.299 0.470 0.509 0.184 0.307 0.226 0.344 0.176 0.315 0.351 0.459 0.219 0.335 0.300 0.369
336 0.339 0.424 0.371 0.439 0.331 0.417 0.351 0.432 0.301 0.397 1.268 0.883 0.349 0.431 0.367 0.448 0.313 0.427 1.324 0.853 0.421 0.476 0.509 0.524
720 0.923 0.726 0.900 0.714 0.847 0.691 0.886 0.714 0.901 0.714 1.767 1.068 0.852 0.698 0.964 0.746 0.839 0.695 1.058 0.797 1.092 0.769 1.447 0.941
AVG 0.380 0.413 0.417 0.417 0.360 0.403 0.378 0.417 0.367 0.404 0.940 0.707 0.370 0.413 0.416 0.443 0.354 0.414 0.750 0.626 0.461 0.454 0.613 0.539
$1^{\mathrm{st}}$ Count 14 27 13 0 0 3 0 0 1 1 0 0 0 0 0 0 2 0 0 0 0 0 0 0

### v4

@@
-ProtoTS: Learning Hierarchical Prototypes for Explainable Time Series Forecasting
+arXiv:2509.23159v4[cs.LG] 27 Feb 2026
 Abstract
@@
 Experiments on multiple realistic benchmarks, including a newly released LOF dataset, show that ProtoTS not only exceeds existing methods in forecast accuracy but also delivers expert-steerable interpretations for better model understanding and decision support.
+The source code is available at https://github.com/SKURA502/ProtoTS.
 1 Introduction
-Time series forecasting has been widely applied in high-stakes scenarios such as load forecasting (jiang2024interpretable; yang2023interactive), energy management (deb2017review; weron2014electricity), weather prediction (angryk2020multivariate; karevan2020transductive), all of which involve considerable financial impacts.
+Time series forecasting has been widely applied in high-stakes scenarios such as load forecasting (Jiang et al., 2024; Yang et al., 2023), energy management (Deb et al., 2017; Weron, 2014), weather prediction (Angryk et al., 2020; Karevan and Suykens, 2020), all of which involve considerable financial impacts.
 In these applications, while achieving high forecast accuracy is crucial, understanding why and how the model makes specific predictions is equally important.
-It aids in preventing substantial financial losses and building the trust necessary (rojat2021explainable).
-A range of explainable time series forecasting methods have been developed to simultaneously ensure interpretability and good predictive performance (oreshkin2019n; lim2021temporal; zhao2024disentangled; lin2024cyclenet).
+It aids in preventing substantial financial losses and building the trust necessary (Rojat et al., 2021).
+A range of explainable time series forecasting methods have been developed to simultaneously ensure interpretability and good predictive performance (Oreshkin et al., 2019; Lim et al., 2021; Zhao et al., 2024; Lin et al., 2024).
 However, their overall interpretability and potential for further performance improvement are limited, since they mainly provide local, partial explanations for both the output and input sides:
-  • C1: For the output side, existing methods (lim2021temporal; zhao2024disentangled) mainly explain the prediction at individual time steps, lacking the ability to help users quickly interpret the reasons behind the overall trend in the forecast curve.
+  • C1: For the output side, existing methods (Lim et al., 2021; Zhao et al., 2024) mainly explain the prediction at individual time steps, lacking the ability to help users quickly interpret the reasons behind the overall trend in the forecast curve.
 For example, they fail to explain why three peaks occur in the predicted electricity load curve at the noon, afternoon, and night, respectively, and why the three peaks gradually descend (Figure 1(a1)).
 Understanding such overall temporal patterns is of great importance in real-world scenarios.
-For example, power system experts need to ensure demand peaks are identified correctly, so that they can make dispatch decisions (6721569), such as whether additional electricity should be purchased from external sources.
+For example, power system experts need to ensure demand peaks are identified correctly, so that they can make dispatch decisions (Thanos et al., 2013), such as whether additional electricity should be purchased from external sources.
 Failing to provide such analyses limits interpretability and prevents experts from steering the model to improve its accuracy.
-  • C2: For the input side, existing explanations are often limited to certain types of variables, such as focusing solely on endogenous variables (lin2024cyclenet).
+  • C2: For the input side, existing explanations are often limited to certain types of variables, such as focusing solely on endogenous variables (Lin et al., 2024).
 However, as shown in Figure 1(a), there are many different types of input variables, and understanding how their interactions impact the temporal patterns is critical.
@@
 Time Series Forecasting with Exogenous Variables Time series forecasting with exogenous variables has been extensively explored in both classical and modern approaches.
-Traditional statistical models such as ARIMAX (williams2001multivariate) and SARIMAX (vagropoulos2016comparison) extend the ARIMA framework to incorporate correlations between exogenous and endogenous variables.
-Seeking richer dynamics, Transformer families like TFT (lim2021temporal) and TimeXer (wang2024timexer), introduce attention mechanisms to tackle this task.
-Parallel work in MLPs pursues efficiency and long-horizon stability: NBEATSx (oreshkin2019n) appends a dedicated stack for auxiliary factors on top of a residual basis expansion, TiDE (das2023long) factorizes temporal and feature projections with two dense encoders to harness known‑future variables, and TSMixer (ekambaram2023tsmixer; chen2023tsmixer) alternates mixing layers to effectively model their interactions.
-Outside deep learning, generic tabular learners such as XGBoost (chen2016xgboost), LightGBM (ke2017lightgbm), and GAM (yang2023interactive), perform well when exogenous variables are informative and strongly correlated with the target.
+Traditional statistical models such as ARIMAX (Williams, 2001) and SARIMAX (Vagropoulos et al., 2016) extend the ARIMA framework to incorporate correlations between exogenous and endogenous variables.
+Seeking richer dynamics, Transformer families like TFT (Lim et al., 2021) and TimeXer (Wang et al., 2024), introduce attention mechanisms to tackle this task.
+Parallel work in MLPs pursues efficiency and long-horizon stability: NBEATSx (Oreshkin et al., 2019) appends a dedicated stack for auxiliary factors on top of a residual basis expansion, TiDE (Das et al., 2023) factorizes temporal and feature projections with two dense encoders to harness known‑future variables, and TSMixer (Ekambaram et al., 2023; Chen et al., 2023) alternates mixing layers to effectively model their interactions.
+Outside deep learning, generic tabular learners such as XGBoost (Chen and Guestrin, 2016), LightGBM (Ke et al., 2017), and GAM (Yang et al., 2023), perform well when exogenous variables are informative and strongly correlated with the target.
 Our method, ProtoTS, achieves strong performance on this task while offering interpretability and steerability for domain experts, effectively handling large amounts of heterogeneous variables while filtering out irrelevant information.
 Time Series Interpretability Time series interpretability can be divided into post-hoc and ante-hoc approaches.
-Post-hoc interpretability focuses on unveiling the reasoning of pre-existing black-box models (gu2025explainable).
-Generic explanation techniques such as SHAP (lundberg2017unified), LIME (ribeiro2016should), IG (sundararajan2017axiomatic) can be directly applied to time-series models.
-crabbe2021explaining learns sparse perturbation masks over the time-feature dimensions, struggling to quantify each input’s effect on a regression forecast.
+Post-hoc interpretability focuses on unveiling the reasoning of pre-existing black-box models (Gu et al., 2025).
+Generic explanation techniques such as SHAP (Lundberg and Lee, 2017), LIME (Ribeiro et al., 2016), IG (Sundararajan et al., 2017) can be directly applied to time-series models.
+Crabbé and Van Der Schaar (2021) learns sparse perturbation masks over the time-feature dimensions, struggling to quantify each input’s effect on a regression forecast.
 Consequently, regression tasks tend to favour ante-hoc models whose interpretability is built in by design.
-Following the attribution-centric view of post‑hoc explanations, TFT (lim2021temporal) provides explorable attention over input time steps, and DiPE-Linear (zhao2024disentangled) parameterizes filters in both temporal and spectral domains to visualise the regions of interest.
-N-BEATS (oreshkin2019n) uses residual blocks for generic, trend, and seasonal components, and olivares2023neural extends the scheme to exogenous variables.
+Following the attribution-centric view of post‑hoc explanations, TFT (Lim et al., 2021) provides explorable attention over input time steps, and DiPE-Linear (Zhao et al., 2024) parameterizes filters in both temporal and spectral domains to visualise the regions of interest.
+N-BEATS (Oreshkin et al., 2019) uses residual blocks for generic, trend, and seasonal components, and Olivares et al.
+(2023) extends the scheme to exogenous variables.
 These works are limited to local explanations and fail to provide a global view that experts need for a shared understanding.
-CycleNet (lin2024cyclenet) identifies common periodic patterns in time-series data but only focuses on endogenous variables.
+CycleNet (Lin et al., 2024) identifies common periodic patterns in time-series data but only focuses on endogenous variables.
 In contrast, ProtoTS holds a global set of prototypical patterns and explains each forecast by matching it to them, offering both global overview and local interpretation.
@@
 These methods typically fall into two categories.
-1) prototypes decoded as intermediate representations (shen2025learn; li2023prototype; jin2023time) or model parameters (chen2024similarity).
-For example, jin2023time uses prototypes to map time-series embeddings into textual embeddings for large language models.
+1) prototypes decoded as intermediate representations (Shen, 2025; Li et al., 2023a; Jin et al., 2023a) or model parameters (Chen et al., 2024).
+For example, Jin et al.
+(2023a) uses prototypes to map time-series embeddings into textual embeddings for large language models.
 Such designs enrich the model’s architecture but leave the prototypes detached from the output.
-2) prototypes directly mapped to output variables (queen2023encoding; liu2024timex++; ming2019interpretable; obermair2023example), serving as typical examples of the outputs.
+2) prototypes directly mapped to output variables (Queen et al., 2023; Liu et al., 2024; Ming et al., 2019; Obermair et al., 2023), serving as typical examples of the outputs.
 While the latter offers stronger interpretability, existing approaches decode a prototype into only a single output variable (e.g., a class prediction).
@@
 Formally, let the historical endogenous variables (e.g., electric load) as $\mathbf{Y}_{1:L}=\{\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{L}\}\in\mathbb{R}^{L\times 1}$ and the associated exogenous variables (e.g., temperature and holiday) as $\mathbf{X}_{1:L}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{L}\}\in\mathbb{R}^{L\times C}$, where $L$ is the look-back window size and $C$ is the number of exogenous variables.
-In many practical scenarios (wessel2020using; lin2023prediction), future exogenous variables $\mathbf{X}_{L+1:L+H}=\{\mathbf{x}_{L+1},\dots,\mathbf{x}_{L+H}\}\in\mathbb{R}^{H\times C}$ are also available.
+In many practical scenarios (Wessel, 2020; Lin et al., 2023), future exogenous variables $\mathbf{X}_{L+1:L+H}=\{\mathbf{x}_{L+1},\dots,\mathbf{x}_{L+H}\}\in\mathbb{R}^{H\times C}$ are also available.
 For example, in the LOF dataset, we obtain forecasted weather variables (e.g., temperature) from commercial weather services, and calendar variables (e.g., holiday flags) are accessible in advance.
@@
 Multi-Channel Embedding In ProtoTS, we adopt a multi-channel embedding where endogenous and exogenous variables are processed separately to achieve optimal encoding for each type.
-Specifically, endogenous variables are directly encoded through an endogenous encoding channel, implemented as a non-linear projection $\gamma$ using a multi-layer perceptron with activation functions, a widely used design (das2023long).
-For exogenous variables, we follow standard practice in tabular learning (wang2022transtab).
+Specifically, endogenous variables are directly encoded through an endogenous encoding channel, implemented as a non-linear projection $\gamma$ using a multi-layer perceptron with activation functions, a widely used design (Das et al., 2023).
+For exogenous variables, we follow standard practice in tabular learning (Wang and Sun, 2022).
 We categorize them into discrete variables (e.g., is holiday, day of week) and continuous variables (e.g., temperature, humidity), and apply tailored embedding methods to best capture the properties of each variable.
@@
 Continuous variables $\mathbf{x}^{\text{con}}_{t,j}$ are projected into the embedding space through a continuous encoding channel, applying variable-specific non-linear projections $\psi_{j}$.
-This multi-channel embedding allows the model to better preserve the information contained in heterogeneous variables.
-Bottleneck Channel Fusion To effectively fuse all variables information at time step $t$ and obtain a full representation, we aggregate the embeddings of all variables through addition, which integrates information across heterogeneous variables without additional parameters (arora2017simple):
+This multi-channel embedding allows the model to better preserve the information contained in heterogeneous variables Munz-Körner and Weiskopf (2024).
+Bottleneck Channel Fusion To effectively fuse all variables information at time step $t$ and obtain a full representation, we aggregate the embeddings of all variables through addition, which integrates information across heterogeneous variables without additional parameters (Arora et al., 2017):
 $\mathbf{Z}_{t}=\begin{cases}\gamma(\mathbf{y}_{t})+\sum\limits_{j=1}^{C_{\text{dis}}}\mathbf{E}_{j}(\mathbf{x}^{\text{dis}}_{t,j})+\sum\limits_{j=1}^{C_{\text{con}}}\psi_{j}(\mathbf{x}^{\text{con}}_{t,j}),&t\in[1:L]\quad\text{ (Look-back Window)}\\ \sum\limits_{j=1}^{C_{\text{dis}}}\mathbf{E}_{j}(\mathbf{x}^{\text{dis}}_{t,j})+\sum\limits_{j=1}^{C_{\text{con}}}\psi_{j}(\mathbf{x}^{\text{con}}_{t,j}),&t\in[L+1:L+H]\quad\text{ (Forecast Window)}\end{cases}$ (2)
@@
 Concretely, we project the high-dimensional embeddings into a lower-dimensional latent space through a bottleneck projection, $\mathbb{R}^{d}\to\mathbb{R}^{d_{\text{bottle}}}\to\mathbb{R}^{d}$, where $d_{\text{bottle}}\ll d$.
-The compressed representation is then mapped back to the original space, retaining the most relevant information while filtering out noise (wang2021variational).
-We apply the bottleneck layer within an MLP-Mixer architecture (tolstikhin2021mlp), inserting it into both $\text{MLP}_{\text{feature}}$ and $\text{MLP}_{\text{time}}$.
+The compressed representation is then mapped back to the original space, retaining the most relevant information while filtering out noise (Wang et al., 2021).
+We apply the bottleneck layer within an MLP-Mixer architecture (Tolstikhin et al., 2021), inserting it into both $\text{MLP}_{\text{feature}}$ and $\text{MLP}_{\text{time}}$.
 Fusion is performed sequentially, first along the feature dimension and then along the temporal dimension, T denotes the transpose of the vector:
@@
 where $\text{W}\in\mathbb{R}^{(L+H)\times 1}$ projects the sequence length from $L+H$ to 1.
-Prototype Similarity Computing To provide a user-friendly global explanation while handling complex information at a local level for detailed interpretation and accurate forecasting, ProtoTS integrates prototypical explanation directly into its prediction mechanism via a learnable set of prototypes $\mathbf{\Pi}$.
+Prototype Similarity Computing To provide a user-friendly global explanation while handling complex information at a local level for detailed interpretation and accurate forecasting, ProtoTS integrates prototypical explanation directly into its prediction mechanism via a learnable set of prototypes $\mathbf{\Pi}$ (Zhang et al., 2024b).
 Each prototype in $\mathbf{\Pi}$ is represented as an embedding $\bm{\mu}\in\mathbb{R}^{d}$ and its corresponding temporal pattern $\mathbf{p}\in\mathbb{R}^{T}$, which are both learnable parameters.
@@
 The distance can be chosen flexibly, such as Euclidean distance, cosine similarity.
-We adopt Euclidean distance, under which the prototypes are equivalent to a Gaussian mixture model in the embedding with an identity covariance matrix (allen2019infinite).
+We adopt Euclidean distance, under which the prototypes are equivalent to a Gaussian mixture model in the embedding with an identity covariance matrix (Allen et al., 2019).
 Meanwhile, temporal pattern $\mathbf{p}$ is used to form the final prediction.
-Specifically, we fix each prototype temporal pattern $\mathbf{p}$ over a predefined period of length $T$, and during forecasting, the prototype pattern is aligned with the forecast window $H$ based on its phase and length (lin2024cyclenet).
+Specifically, we fix each prototype temporal pattern $\mathbf{p}$ over a predefined period of length $T$, and during forecasting, the prototype pattern is aligned with the forecast window $H$ based on its phase and length (Lin et al., 2024).
 3.3 Hierarchical Prototype Learning
@@
 $\widehat{\mathbf{Y}}_{L+1:L+H}=\sum_{i=1}^{N}f(\hat{\mathbf{Z}}|\bm{\mu}_{i})\sum_{j=1}^{M}f(\hat{\mathbf{Z}}|\bm{\mu}_{i,j})\cdot\mathbf{p}_{i,j}.$ (8)
-While the above describes a two-level hierarchy, the structure is general: any leaf prototype can be incrementally split to form a multi-level tree.
-However, not all prototypes need splitting, and some splits may yield overly similar temporal patterns.
-Therfore, we propose a splitting rule for determining which leaf prototypes should be further refined.
-The rule is designed to identify prototypes whose current temporal patterns are insufficiently representative of their associated instances and thus require additional fine-grained refinement.
-The pseudocode is provided in Algorithm 1.
 Algorithm 1 Prototype Splitting Rule
@@
 Select leaf prototypes with top $\alpha\%$ values in $\text{NormLoss}[l]$.
+While the above describes a two-level hierarchy, the structure is general: any leaf prototype can be incrementally split to form a multi-level tree.
+However, not all prototypes need splitting, and some splits may yield overly similar temporal patterns.
+Therefore, we propose a splitting rule for determining which leaf prototypes should be further refined.
+The rule is designed to identify prototypes whose current temporal patterns are insufficiently representative of their associated instances and thus require additional fine-grained refinement.
+The pseudocode is provided in Algorithm 1.
 This hierarchical design enables experts to first obtain a global understanding through high-level root prototypes (e.g., distinguishing between seasonal or holiday-related patterns), and then refine their insights by examining lower-level child prototypes that capture more localized variations (e.g., differentiating between long holidays and short holidays).
@@
 Datasets Load Forecasting dataset (LOF) is a synthetic electric load dataset, including 22 supporting covariates and covering four regions.
-We also evaluated on the Electricity Price Forecasting (EPF) dataset (lago2021forecasting), which covers hourly electric prices from five Nordic markets.
+We also evaluated on the Electricity Price Forecasting (EPF) dataset (Lago et al., 2021), which covers hourly electric prices from five Nordic markets.
 The specific details of the datasets can be found in the Appendix B.1.
-Baselines To provide a wide‑ranging comparison, we benchmark eight models in total: five state-of-the-art deep-learning methods (TimeXer (wang2024timexer), iTransformer (liu2023itransformer), TiDE (das2023long), TFT (lim2021temporal), N‑BEATSx (olivares2023neural)) together with three strong machine-learning baselines (XGBoost (chen2016xgboost), LightGBM (ke2017lightgbm), pyGAM (serven2018pygam)).
+Baselines To provide a wide‑ranging comparison, we benchmark eight models in total: five state-of-the-art deep-learning methods (TimeXer (Wang et al., 2024), iTransformer (Liu et al., 2023), TiDE (Das et al., 2023), TFT (Lim et al., 2021), N‑BEATSx (Olivares et al., 2023)) together with three strong machine-learning baselines (XGBoost (Chen and Guestrin, 2016), LightGBM (Ke et al., 2017), pyGAM (Servén and Brummitt, 2018)).
 All deep-learning models support exogenous variables as inputs, and we further extend TimeXer and iTransformer to incorporate future covariates.
@@
 Visual comparison can be found in Appendix E.
+We also evaluate ProtoTS on the EPF dataset.
+Unlike LOF, this task involves fewer covariates but exhibits higher volatility, posing challenges for accurate forecasting.
+As shown in Table 2, ProtoTS achieves the best performance across all five markets, reducing MSE by 8% and MAE by 8% compared to the best-performing baseline.
+We further conduct a sensitivity analysis under 5 random seeds, see Appendix C.3.
+Full implementation details can be found in the Appendix B.2.
 Table 1: MAE results on the LOF dataset with 22 supporting covariates.
@@
 $\Delta$ - 32% 29% 21% 50% 81% 48% 44% 52%
-We also evaluate ProtoTS on the EPF dataset.
-Unlike LOF, this task involves fewer covariates but exhibits higher volatility, posing challenges for accurate forecasting.
-As shown in Table 2, ProtoTS achieves the best performance across all five markets, reducing MSE by 8% and MAE by 8% compared to the best-performing baseline.
-We further conduct a sensitivity analysis under 5 random seeds, see Appendix C.3.
-Full implementation details can be found in the Appendix B.2.
 Table 2: MAE results on EPF dataset equipped with 6 supporting covariates.
@@
 As the available data decreased, existing baselines such as TimeXer, iTransformer, and TiDE show clear drops in MSE.
-In contrast, ProtoTS maintains stable results with only minor performance loss, which improves high data efficiency.
+In contrast, ProtoTS maintains stable results with only minor performance loss, which improves high data efficiency (Jin et al., 2023b).
 The comparison results are illustrated in Figure 3(b).
@@
 Users were asked concrete questions about the input variables, such as "Given this prediction and explanations, which season does the predicted day most likely belong to?".
-To evaluate interpretability, we define User Precision (UPrec) as the proportion of correctly answered variable-related questions, reflecting whether users can accurately infer the values of relevant input features from the explanations.
-To further assess usability, we adopt the System Usability Scale (SUS) (brooke1996sus), a widely used questionnaire that measures users’ perceived ease of use and overall satisfaction with a system.
+To evaluate interpretability, we define User Precision (UPrec) as the proportion of correctly answered variable-related questions, reflecting whether users can accurately infer the values of relevant input features from the explanations (Lee et al., 2025).
+To further assess usability, we adopt the System Usability Scale (SUS) (Brooke and others, 1996), a widely used questionnaire that measures users’ perceived ease of use and overall satisfaction with a system.
 The quantitative results are reported in Table 4, with additional details in Appendix D.
@@
 Extensive experiments and case studies on realistic datasets show state-of-the-art performance and transparent support, enabling experts to refine model behavior and bridging accuracy with interpretability.
-Ethics statement
-We provide an ethics statement as follows:
-  • Human subjects research: Our experiments include a user study (see Section 4.3). The study does not raise any ethical concerns, and all participants gave informed consent prior to participation.
-  • Datasets: All datasets used in our work are publicly available. The released dataset is a simulation of real-world data and does not contain any actual sensitive information.
-  • Fairness and bias: Our work does not raise fairness or bias concerns.
-Reproducibility statement
-We confirm that the contents of the main text and the appendix are sufficient to reproduce our work and our experiments are reproducible.
-Section 3 clearly describes the proposed method and workflow.
-Appendix B.1 provides detailed dataset descriptions, and Appendix B.2 provides sufficient implementation details (including hyperparameters and training settings).
-We also include in the abstract a link to an anonymous repository containing our source code and the released dataset; the repository provides scripts for data preparation, training, and evaluation, along with configuration files and random seeds.
-We have verified that the code is consistent with the descriptions in the paper and that the reported experiments can be reproduced following the provided instructions.
+Acknowledgements
+This work was partly supported by the National Natural Science Foundation of China (NSFC) (NO.
+62476279, NO.
+U2436209), Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, Renmin University of China, the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China No.24XNKJ18.
+This work was partially supported by fund for building world-class universities (disciplines) of Renmin University of China and Public Computing Cloud, Renmin University of China.
+This work was partially supported by DAMO Academy (Hupan Laboratory) through DAMO Academy (Hupan Laboratory) Innovative Research Program.
 Appendix A Use of LLMs
@@
 LOF dataset The LOF dataset is an synthetic electric load forecasting dataset.
-It is designed to predict future electricity demand at the provincial and city levels, enable electricity generators, distributors, and suppliers to plan effectively ahead and promote energy conservation among the users (nti2020electricity).
+It is designed to predict future electricity demand at the provincial and city levels, enable electricity generators, distributors, and suppliers to plan effectively ahead and promote energy conservation among the users (Nti et al., 2020).
 Due to the influence of human industrial and commercial activities, electricity demand exhibits a clear daily cycle, maintaining a relatively stable and predictable pattern under normal conditions.
@@
 The specific details of these exogenous variables are presented in Table 10.
-EPF dataset The EPF dataset is an electricity price forecasting datasets (lago2021forecasting).
+EPF dataset The EPF dataset is an electricity price forecasting datasets (Lago et al., 2021).
 It collects from five distinct day-ahead electricity markets in Northern Europe, covering a six year period with a sampling interval of 1 hour.
@@
 B.2 Implementation Details
-All experiments are implemented using PyTorch (paszke2019pytorch) and executed on a single NVIDIA V100 GPU with 32GB of memory.
-For model optimization, we employ the Adam (kingma2014adam) optimizer with an initial learning rate of and use our Loss as the objective function.
+All experiments are implemented using PyTorch (Paszke, 2019) and executed on a single NVIDIA V100 GPU with 32GB of memory.
+For model optimization, we employ the Adam (Kingma, 2014) optimizer with an initial learning rate of and use our Loss as the objective function.
 The training is set for a maximum of 30 epochs, with early stopping applied to prevent overfitting.
@@
 All baseline models used for comparison are re-implemented based on the TimeXer Repository benchmark.
-All baseline use REVIN (kim2021reversible) in training and test process.
+All baseline use REVIN (Kim et al., 2021) in training and test process.
 Appendix C Full Results
@@
 The second part is System Usability Scale (SUS) Questionnaire, designed to assess the usability of the explanation systems offering different levels of explainability.
-Participants are asked to subjectively rate the system based on their experience completing the first part.
+Participants are asked to subjectively rate the system based on their experience completing the first part (Xu et al., 2024).
 The SUS yields a single number representing a composite measure of the overall usability of the system under study.
@@
 However, this process requires sufficient domain knowledge and manual effort.
-In domains where expert resources are hard to access or knowledge is difficult to formalize, the usability and effectiveness of this feature may be limited.
+In domains where expert resources are hard to access or knowledge is difficult to formalize, the usability and effectiveness of this feature may be limited (Zhang et al., 2024a).
   • Computational cost: ProtoTS introduces additional computational overhead compared to transformer-based or MLP-based baselines (e.g., iTrans and TiDE).
@@
 Datasets We conduct comprehensive experiments on six benchmark datasets to evaluate the multivariate forecasting capability of ProtoTS. These datasets cover a variety of domains:
-  • ETTh1 and ETTh2 (zhou2021informer): Hourly electricity load data from two different nodes in a power grid.
+  • ETTh1 and ETTh2 (Zhou et al., 2021): Hourly electricity load data from two different nodes in a power grid.
 They capture seasonal patterns and variable interactions across multiple features, serving as standard benchmarks for long-term forecasting.
-  • ETTm1 and ETTm2 (zhou2021informer): Minute-level electricity load datasets focusing on finer temporal granularity.
+  • ETTm1 and ETTm2 (Zhou et al., 2021): Minute-level electricity load datasets focusing on finer temporal granularity.
 They emphasize short-term dynamics and present challenges in modeling high-frequency fluctuations.
-  • Weather (zhou2021informer): A multivariate weather dataset containing temperature, humidity, wind speed, and other meteorological variables collected from multiple stations.
-  • Exchange (lai2018modeling): A financial time series dataset with daily exchange rates of eight foreign currencies.
+  • Weather (Zhou et al., 2021): A multivariate weather dataset containing temperature, humidity, wind speed, and other meteorological variables collected from multiple stations.
+  • Exchange (Lai et al., 2018): A financial time series dataset with daily exchange rates of eight foreign currencies.
 For all datasets, we follow standard experimental protocols, using a 96 look-back window and forecasting multiple future steps {96, 192, 336, 720} to evaluate both short-term accuracy and long-horizon stability.
-Baseline We compare ProtoTS against a broad set of baselines, including: Transformer-based models (TimeXer (wang2024timexer), iTransformer (liu2023itransformer), PatchTST (nie2022time), TimesNet (wu2022timesnet), Crossformer (zhang2023crossformer), Stationary (liu2022non), Autoformer (wu2021autoformer)), MLP-based architectures (TiDE (das2023long), DLinear (zeng2023transformers), RLinear (li2023revisiting), SCINet (liu2022scinet)).
+Baseline We compare ProtoTS against a broad set of baselines, including: Transformer-based models (TimeXer (Wang et al., 2024), iTransformer (Liu et al., 2023), PatchTST (Nie et al., 2022), TimesNet (Wu et al., 2022), Crossformer (Zhang and Yan, 2023), Stationary (Liu et al., 2022b), Autoformer (Wu et al., 2021)), MLP-based architectures (TiDE (Das et al., 2023), DLinear (Zeng et al., 2023), RLinear (Li et al., 2023b), SCINet (Liu et al., 2022a)).
 Results The results in Table 11 indicate that ProtoTS achieves leading performance across most multivariate forecasting tasks.
@@
 Table 11: Full results of multivariate time series forecasting task, where the input window is set to 96, and the forecast window is set to {96, 192, 336, 720}.
-Results are citepd from TimeXer (wang2024timexer), otherwise reproduced.
+Results are citepd from TimeXer (Wang et al., 2024), otherwise reproduced.
 $1^{\mathrm{st}}$ denotes the first count.
@@
 Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-Etth1
-96 0.380 0.388 0.382 0.403 0.386 0.405 0.386 0.395 0.414 0.419 0.423 0.448 0.479 0.464 0.384 0.402 0.386 0.400 0.654 0.599 0.513 0.491 0.449 0.459
+Etth1 96 0.380 0.388 0.382 0.403 0.386 0.405 0.386 0.395 0.414 0.419 0.423 0.448 0.479 0.464 0.384 0.402 0.386 0.400 0.654 0.599 0.513 0.491 0.449 0.459
 192 0.437 0.423 0.429 0.435 0.441 0.436 0.437 0.424 0.460 0.445 0.471 0.474 0.525 0.492 0.436 0.429 0.437 0.432 0.719 0.631 0.534 0.504 0.500 0.482
@@
 AVG 0.445 0.427 0.437 0.437 0.454 0.447 0.446 0.434 0.469 0.454 0.529 0.522 0.541 0.507 0.458 0.450 0.456 0.452 0.747 0.647 0.570 0.537 0.496 0.487
-Etth2
-96 0.284 0.337 0.286 0.338 0.297 0.349 0.288 0.338 0.302 0.348 0.745 0.584 0.400 0.440 0.340 0.374 0.333 0.387 0.707 0.621 0.476 0.458 0.346 0.388
+Etth2 96 0.284 0.337 0.286 0.338 0.297 0.349 0.288 0.338 0.302 0.348 0.745 0.584 0.400 0.440 0.340 0.374 0.333 0.387 0.707 0.621 0.476 0.458 0.346 0.388
 192 0.370 0.387 0.363 0.389 0.380 0.400 0.374 0.390 0.388 0.400 0.877 0.656 0.528 0.509 0.402 0.414 0.477 0.476 0.860 0.689 0.512 0.493 0.456 0.452
@@
 AVG 0.366 0.394 0.367 0.396 0.383 0.407 0.374 0.398 0.387 0.407 0.942 0.684 0.611 0.550 0.414 0.427 0.559 0.515 0.954 0.723 0.526 0.516 0.450 0.459
-Ettm1
-96 0.324 0.347 0.318 0.356 0.334 0.368 0.355 0.376 0.329 0.367 0.404 0.426 0.364 0.387 0.338 0.375 0.345 0.372 0.418 0.438 0.386 0.398 0.505 0.475
+Ettm1 96 0.324 0.347 0.318 0.356 0.334 0.368 0.355 0.376 0.329 0.367 0.404 0.426 0.364 0.387 0.338 0.375 0.345 0.372 0.418 0.438 0.386 0.398 0.505 0.475
 192 0.375 0.369 0.362 0.383 0.387 0.391 0.391 0.392 0.367 0.385 0.450 0.451 0.398 0.404 0.374 0.387 0.380 0.389 0.426 0.441 0.459 0.444 0.553 0.496
@@
 AVG 0.392 0.387 0.382 0.397 0.407 0.410 0.414 0.407 0.387 0.400 0.513 0.496 0.419 0.419 0.400 0.406 0.403 0.407 0.485 0.481 0.481 0.456 0.588 0.517
-Ettm2
-96 0.176 0.254 0.171 0.256 0.180 0.264 0.182 0.265 0.175 0.259 0.287 0.366 0.207 0.305 0.187 0.267 0.193 0.292 0.286 0.377 0.192 0.274 0.255 0.339
+Ettm2 96 0.176 0.254 0.171 0.256 0.180 0.264 0.182 0.265 0.175 0.259 0.287 0.366 0.207 0.305 0.187 0.267 0.193 0.292 0.286 0.377 0.192 0.274 0.255 0.339
 192 0.237 0.295 0.237 0.299 0.250 0.309 0.246 0.304 0.241 0.302 0.414 0.492 0.290 0.364 0.249 0.309 0.284 0.362 0.399 0.445 0.280 0.339 0.281 0.340
@@
 AVG 0.275 0.312 0.274 0.322 0.288 0.332 0.286 0.327 0.281 0.326 0.757 0.610 0.358 0.404 0.291 0.333 0.350 0.401 0.571 0.537 0.306 0.347 0.327 0.371
-weather
-96 0.155 0.197 0.157 0.205 0.174 0.214 0.192 0.232 0.177 0.218 0.158 0.230 0.202 0.261 0.172 0.220 0.196 0.255 0.221 0.306 0.173 0.223 0.266 0.336
+weather 96 0.155 0.197 0.157 0.205 0.174 0.214 0.192 0.232 0.177 0.218 0.158 0.230 0.202 0.261 0.172 0.220 0.196 0.255 0.221 0.306 0.173 0.223 0.266 0.336
 192 0.202 0.241 0.204 0.247 0.221 0.254 0.240 0.271 0.225 0.259 0.206 0.277 0.242 0.298 0.219 0.261 0.237 0.296 0.261 0.340 0.245 0.285 0.307 0.367
@@
 AVG 0.238 0.264 0.241 0.271 0.258 0.279 0.272 0.291 0.259 0.281 0.259 0.315 0.271 0.320 0.259 0.287 0.265 0.317 0.292 0.363 0.288 0.314 0.338 0.382
-Exchange
-96 0.085 0.205 0.090 0.209 0.086 0.206 0.093 0.217 0.088 0.205 0.256 0.367 0.094 0.218 0.107 0.234 0.088 0.218 0.267 0.396 0.111 0.237 0.197 0.323
+Exchange 96 0.085 0.205 0.090 0.209 0.086 0.206 0.093 0.217 0.088 0.205 0.256 0.367 0.094 0.218 0.107 0.234 0.088 0.218 0.267 0.396 0.111 0.237 0.197 0.323
 192 0.173 0.299 0.190 0.309 0.177 0.299 0.184 0.307 0.176 0.299 0.470 0.509 0.184 0.307 0.226 0.344 0.176 0.315 0.351 0.459 0.219 0.335 0.300 0.369